OpenStudy (anonymous):

jason considered two simialar televisons at the local store. the genetic version was based off the brand name and was 2/3 the size of the brand name .if the genetic television set is 12 inches by 24 inches , what are the dimensions of the brand name television?

OpenStudy (anonymous):

The generic TV is 3/4 as big as the brand name TV. If 'l' and 'w' are the length & width of the brand name TV, then... 24 = (3/4)l and 12 = (3/4)w (4/3)24 = l and (4/3)12 = w 32 = l and 16 = w The brand name TV is 32 x 16.

OpenStudy (anonymous):

THANKS ! i didnt get the question at all!